Product Cost Concept of Product Costing

MyPhone Inc. uses the product cost concept of applying the cost-plus approach to product pricing. The costs of producing and selling 5,000 cellular phones are as follows:

Variable costs: Fixed costs:
Direct materials $89 per unit Factory overhead $199,600
Direct labor 37 Selling and administrative expenses 70,000
Factory overhead 24
Selling and administrative expenses 23
Total $173 per unit

MyPhone wants a profit equal to a 15% rate of return on invested assets of $601,400.

a. Determine the amount of desired profit from the production and sale of 5,000 cellular phones.
$

b. Determine the product cost and the cost amount per unit for the production of 5,000 cellular phones. If required, round your answer to nearest dollar.
$ per unit

c. Determine the product cost markup percentage (rounded to two decimal places) for cellular phones.
%

d. Determine the selling price of cellular phones. Round to the nearest dollar.

Cost $ per unit
Markup
Selling price $ per unit
Answer #1

Solution a:

Desired profit = Invested assets *15% = $601400 *15% = $90,210

Solution b:

Total product Cost = Variable + Fixed = [5000*($89+$37+$24)] + $199600 = $750,000 = $949,600

Cost per unit = $949600 / 5000 = $189.92 per unit = 190 per unit

Solution c:

Total markup = Desired profit + Selling and administrative expense = $90210 + (5000*$23 + $70000) = $275,210

Markup percentage = Total markup / Total product cost = $275,210 / $949,600 = 28.98%

Solution d:

Cost = $190

Markup = 190 *28.98% = $55

Selling price = $190 + $55 = $245